owned and managed by Donna Mason The Mason Inn, Inc., a rooms-only property owned and mana itable year in 20X7. Donna has requested your assistance preparing the ans get. Assume the lodging operation is open 365 days during 20x8. She provides with the following information for 20X8:
media%2Fcab%2Fcabb0b63-6faa-4cc8-a18e-fb
Answer #1

Answer:

We have to prepare the annual budgets of 20x8,

Particulars Amount($)
Room sales:(Number*Percentage*ADR*Days)
Single rooms= (30*70%*$80**365) 613,200
Double rooms= (40*60%*$90*365) 788,400
Kings rooms= (30*80%*$100*365) 876,000
Total =613000+788400+876000 2,277,600
Other revenues=2277600*(4/100-4) 94,900
Total revenue=2277600+94900 2,372,500
Room expenses:
Labour=(100,000+15% of 2277600) 441,640
Supplies=(5*(30*70%+40*60%+30*80%)*365) 125,925
Other=(10000+2% of 2277600) 55,552
Total Expenses 623,117
Operating expenses
A & G=(150000+3% of 2277600) 218,328
S & M=(30000+6% of 2277600) 166,656
Franchise fees=(6% of 2277600) 136,656
Maintainence=(60000+4% of 2277600) 151,104
Energy=(30000+4% of 2277600) 1211,04
Total 793,848
Fixed charges
Insurance 20,000
Property tax 40,000
Depreciation 50,000
Total 110,000
Not operating income=2372500-623117-793848-11000) 845,535
interest expenses=8% of 1000000 80,000
Pre tax income=845535-80000 765,535
taxes=20% of 765535 153,107
Not income=765535-153107 612,428